# CHROME: Clothed Human Reconstruction with Occlusion-Resilience and Multiview-Consistency from a Single Image

### v1

@@ 2 Related Works
Clothed Human Reconstruction: Clothed human reconstruction focuses on generating a 3D mesh of clothed subjects, textured or untextured, from various inputs such as single images, multiview images, or monocular videos. Existing methods fall primarily into two categories: implicit function-based techniques [1, 17, 24, 18, 9, 10, 2] and rendering-based approaches [5, 21, 41, 42]. Implicit function-based methods, such as PIFu [1] and its extensions [8], leverage pixel-aligned features processed through learnable decoders to reconstruct detailed 3D models using 3D supervision. Recent approaches like GTA [9] and SIFU [2] incorporate the SMPL model to embed anatomical priors, improving the fidelity of reconstruction. TeCH [43] applies optimization techniques to achieve single-image 3D digitization. However, occlusions persist as a critical challenge in clothed human reconstruction. Wang et al. [44] proposed a methodology utilizing ground-truth SMPL meshes to mitigate occlusion effects. However, their framework presents two principal limitations: heavy reliance on precise SMPL mesh estimations, which are inherently challenging to acquire under occluded conditions, and lack of publicly available codes, obstructing both reproducibility and practical adoption. With advances in Neural Radiance Fields (NeRFs) [19] and 3D Gaussian Splatting (3DGS) [20], studies such as [5, 21] have adapted these techniques for human reconstruction from monocular videos, though occlusion handling remains problematic, as evidenced in occlusion-robust variants [22, 45, 12]. Despite this progress, they typically require monocular video inputs and are optimized per subject, limiting their scalability and practical deployment. Differently, we propose a single image-based generalizable approach for occlusion-free novel view synthesis and geometric reconstruction.
@@
Multi-view Diffusion and Large Reconstruction Models for 3D Reconstruction: Latent Diffusion Models (LDMs) [46] have aided significant advancements in image generation, reconstruction, and single-image-to-3D tasks. Zero-1-to-3 [47] pioneered the use of LDMs to generate novel views from single input images with specified camera parameters, sparking a wave of related research [23, 48, 49, 50]. ControlNet [51] introduced guided control over LDM sampling, advancing conditional generation. Recent works facilitate the creation of 3D objects from text prompts [52, 53] by leveraging extensive training datasets [54]. Pretrained LDMs also function as priors for novel view synthesis, modeling natural image distributions, and penalizing the reconstruction model through Score Distillation Sampling based losses [55, 56, 57]. Recently, the Large Reconstruction Model (LRM) [58] introduced NeRF-based triplane representations for efficient single image to 3D generation, trained on large multiview datasets [59, 60]. Large Gaussian Model (LGM) [61] extended the same by integrating pretrained diffusion models for single-image-to-3D synthesis, achieving faster rendering with 3DGS techniques. However, these algorithms struggle with partial occlusions, motivating our proposed approach for 3D human recovery from single or stereo occluded images.
@@ 3.1 Pose Controlled Multiview Consistent Diffusion
Figure 3: Qualitative analysis of $\mathcal{F}_{D}$ in generating occlusion-free novel views: Qualitative results demonstrating the capacity of $\mathcal{F}_{D}$ in generating occlusion-free novel views conditioned on a single occluded image from CustomHumans (left) and AHP (right).
@@ 3.2 Novel View Synthesis using Gaussian Splatting
Figure 4: Qualitative analysis of CHROME on artificially occluded THuman2.0 and CustomHumans: Qualitative results of CHROME against state-of-the-art algorithms [1, 9, 2, 10] on artificially occluded THuman2.0 (top three rows) and artificially occluded CustomHumans (bottom three rows); clearly none of the baseline algorithm perform occlusion-free novel view synthesis whereas CHROME effectively mitigates occlusions and predicts multiview consistent reconstructions.
@@ 4 Experiments and Results
In this section, we demonstrate CHROME’s outstanding ability to perform occlusion-free novel view synthesis and geometric reconstruction for single-view occluded human image on multiple datasets, presenting both qualitative and quantitative analyses on multiple benchmarks.
Datasets: We evaluate our method on multiple datasets: THuman2.0 [68] (526 high-quality scans of 150 clothing styles across 500 training and 26 evaluation subjects) using a perspective camera (49.1 FOV) with 16 views per subject; CustomHumans [69] containing 600 scans of 80 subjects with 120 garments, augmented with artificial occlusions for zero-shot testing; CAPE [70] containing 15 subjects in tight clothing, modified with synthetic occlusions for zero-shot evaluation and AHP [11] featuring naturally occluded in-the-wild images used for input-view de-occlusion assessment.
Implementation details: For all our experiments, we use the PyTorch coding environment with all models being trained on 4$\times$ A40 GPUs. We train $\mathcal{F_{D}}$ for 100 epochs across all experiments. The training utilizes the AdamW optimizer with a cosine annealing learning rate schedule that peaks at $7\times 10^{-5}$ within 1000 warm-up steps, after which we use a constant learning rate of $5\times 10^{-6}$. During inference, we employ classifier-free guidance with a guidance scale of 4 and 40 diffusion steps, utilizing an ancestral Euler sampling strategy. For training $\mathcal{F_{R}}$, we use a learning rate initialized at $4\times 10^{-4}$, which decays following the Cosine Annealing strategy over 100 epochs. In Equation 4 (see main paper), $\lambda_{1}$ is set to 1.5 and $\lambda_{2}$ is set to 1.
Metrics: We follow [2, 1, 9] and use PSNR, SSIM [71], and LPIPS [67] to evaluate reconstruction accuracy in 2D and chamfer distance (CD) in cm, point-to-surface (P2S) error (cm), and normal consistency (NC) [10] to evaluate 3D reconstruction quality.
@@
Qualitative Results: We present qualitative comparisons of CHROME with the baseline algorithms in Figures 4 and 5 in the in-domain artificially occluded datasets THuman2.0 [68], out-of-domain artificially occluded CustomHumans [69] for novel view synthesis and geometric reconstructions via normal maps. These results clearly show that existing algorithms fail to generate occlusion-free novel views from a single occluded image. Moreover, under occlusions, the generated novel views often suffer from anatomical implausibility and severe multiview inconsistencies. In contrast, CHROME consistently produces novel views that are occlusion-resilient and multiview consistent. This is qualitatively shown for all 16 views in Figure 7. Figure 6 shows the qualitative results of CHROME when stereo inputs are available, showcasing the inherent flexibility of CHROME in generating occlusion-free novel views under stereo reconstruction settings.
Figure 7: Visual Results of CHROME on occluded THuman2.0 and CustomHumans: Qualitative results of CHROME on occluded THuman2.0 and occluded CustomHumans.
Figure 8: Comparative Qualitative Analysis of Existing Large Reconstruction Models for Zero-Shot Novel View Texture Reconstruction from Occluded Single View Images.
How Robust are Existing Large Reconstruction Models for Occlusion-Free Novel View Synthesis? We showcase qualitative comparisons with three baseline algorithms: Zero123++ [23], ImageDream with LGM [61], and SV3D [72], utilizing their pretrained weights for the task of occlusion-free novel view reconstruction in Figure 8. The results reveal significant inconsistencies in these existing methods when applied to our task, underscoring the need for a specialized algorithm CHROME.
Ablation Study: Table 6 highlights the importance of incorporating the estimated 3D pose [35] from the occluded image as an explicit guidance signal for our multiview diffusion model, $\mathcal{F}_{D}$, on the occluded THuman2.0 dataset [68]. In particular, even without using the estimated 3D pose as conditioning information, CHROME outperforms existing baselines as shown in Table 1. However, integrating the pose as explicit guidance leads to more anatomically accurate reconstructions.

### v2

@@ Abstract
$\dagger$ This work was done during Arindam Dutta’s internship at United Imaging Intelligence, Boston, MA.
@@ 2 Related Works
Clothed Human Reconstruction: Clothed human reconstruction focuses on generating a 3D mesh of clothed subjects, textured or untextured, from various inputs such as single images, multiview images, or monocular videos. Existing methods fall primarily into two categories: implicit function-based techniques [1, 17, 24, 18, 9, 10, 2] and rendering-based approaches [5, 21, 41, 42, 43]. Implicit function-based methods, such as PIFu [1] and its extensions [8], leverage pixel-aligned features processed through learnable decoders to reconstruct detailed 3D models using 3D supervision. Recent approaches like GTA [9] and SIFU [2] incorporate the SMPL model to embed anatomical priors, improving the fidelity of reconstruction. TeCH [44] applies optimization techniques to achieve single-image 3D digitization. However, occlusions persist as a critical challenge in clothed human reconstruction [45]. Wang et al. [46] proposed a methodology utilizing ground-truth SMPL meshes to mitigate occlusion effects. However, their reliance on precise SMPL mesh estimations, which are inherently challenging to acquire under occluded conditions obstruct practical adoption. With advances in Neural Radiance Fields (NeRFs) [19] and 3D Gaussian Splatting (3DGS) [20], studies such as [5, 21] have adapted these techniques for human reconstruction from monocular videos, though occlusion handling remains problematic, as evidenced in occlusion-robust variants [22, 47, 12]. Despite this progress, they typically require monocular video inputs and are optimized per subject, limiting their scalability and practical deployment. Differently, we propose a single image-based generalizable approach for occlusion-free novel view synthesis and geometric reconstruction.
@@
Multi-view Diffusion and Large Reconstruction Models for 3D Reconstruction: Latent Diffusion Models (LDMs) [48] have aided significant advancements in image generation, reconstruction, and single-image-to-3D tasks. Zero-1-to-3 [49] pioneered the use of LDMs to generate novel views from single input images with specified camera parameters, sparking a wave of related research [23, 50, 51, 52, 53]. ControlNet [54] introduced guided control over LDM sampling, advancing conditional generation. Recent works facilitate the creation of 3D objects from text prompts [55, 56] by leveraging extensive training datasets [57]. Pretrained LDMs also function as priors for novel view synthesis, modeling natural image distributions, and penalizing the reconstruction model through Score Distillation Sampling based losses [58, 59, 60]. Recently, the Large Reconstruction Model (LRM) [61] introduced NeRF-based triplane representations for efficient single image to 3D generation, trained on large multiview datasets [62, 63]. Large Gaussian Model (LGM) [64] extended the same by integrating pretrained diffusion models for single-image-to-3D synthesis, achieving faster rendering with 3DGS techniques. However, these algorithms struggle with partial occlusions, motivating our proposed approach for 3D human recovery from single or stereo occluded images.
@@ 3.1 Pose Controlled Multiview Consistent Diffusion
Figure 3: Qualitative analysis of $\mathcal{F}_{D}$ in generating occlusion-free novel views: Qualitative results demonstrating the capacity of $\mathcal{F}_{D}$ in generating occlusion-free novel views conditioned on a single occluded image from CustomHumans.
@@ 3.2 Novel View Synthesis using Gaussian Splatting
Figure 4: Qualitative analysis of CHROME on artificially occluded THuman2.0 and CustomHumans: Qualitative results of CHROME against state-of-the-art algorithms [1, 9, 2, 10] on artificially occluded THuman2.0 (top three rows) and artificially occluded CustomHumans (bottom three rows); clearly none of the baseline algorithm perform occlusion-free novel view synthesis whereas CHROME effectively mitigates occlusions and predicts multiview consistent reconstructions. Additional qualitative results are provided in the supplementary.
@@ 4 Experiments and Results
In this section, we demonstrate CHROME’s outstanding ability to perform occlusion-free novel view synthesis and geometric reconstruction for single-view occluded human image on multiple datasets, presenting both qualitative and quantitative analyses on multiple benchmarks. Detailed implementation details and more extensive results e.g. computation time cost, and flexibility with single/stereo-view as input can be found in supplementary material.
Datasets and Metrics: We evaluate our method on multiple datasets: THuman2.0 [71] (526 high-quality scans of 150 clothing styles across 500 training and 26 evaluation subjects) using a perspective camera (49.1 FOV) with 16 views per subject; CustomHumans [72] containing 600 scans of 80 subjects with 120 garments, augmented with artificial occlusions for zero-shot testing; CAPE [73] containing 15 subjects in tight clothing, modified with synthetic occlusions for zero-shot evaluation, AHP [11] featuring naturally occluded in-the-wild images used for input-view de-occlusion assessment and MultiHuman [74] and OCHuman [3] containing multiple people in regular interactions. We follow [2, 1, 9] and use PSNR, SSIM [75], and LPIPS [70] to evaluate reconstruction accuracy in 2D and chamfer distance (CD) in cm, point-to-surface (P2S) error (cm), and normal consistency (NC) [10] to evaluate 3D reconstruction quality.
@@
Figure 6: Clean images: Qualitative Results on occlusion-free images for CHROME against SIFU.
@@
Qualitative Results: We present qualitative comparisons of CHROME with the baseline algorithms in Figures 4 and 5 in the in-domain artificially occluded datasets THuman2.0 [71], out-of-domain artificially occluded CustomHumans [72] for novel view synthesis and geometric reconstructions via normal maps. These results clearly show that existing algorithms fail to generate occlusion-free novel views from a single occluded image. Moreover, under occlusions, the generated novel views often suffer from anatomical implausibility and severe multiview inconsistencies. In contrast, CHROME consistently produces novel views that are resilient to occlusions and consistent across multiple views. Figure 6 shows results of CHROME against SIFU [2] on clean, occlusion-free images wherein clearly CHROME outperforms SIFU. Figure 7 shows the qualitative results of CHROME when stereo inputs are available, showcasing the inherent flexibility of CHROME in generating occlusion-free novel views under stereo reconstruction settings. Additional qualitative results are provided in the supplementary.
Ablation Study: Table 6 highlights the importance of incorporating the estimated 3D pose [35] from the occluded image as an explicit guidance signal for our multiview diffusion model, $\mathcal{F}_{D}$, on the occluded THuman2.0 dataset [71]. In particular, even without using the estimated 3D pose as conditioning information, CHROME outperforms existing baselines as shown in Table 1. However, integrating the pose as explicit guidance leads to more anatomically accurate reconstructions, as shown in Figure 8 and discussed in the supplementary.
@@
Figure 8: Importance of pose conditioning: An example case where explicit pose guidance enhances anatomical accuracy and visual quality in novel view reconstructions. Left: input occluded image. Middle: novel view synthesis by CHROME without pose conditioning. Right: Novel view synthesis by CHROME.
@@ 5 Conclusion
Acknowledgments. This work was partially supported by NSF grants CMMI-2326309 and CNS-2312395.